On a map, the distance between Mary's house and John's house is 9.25 inches. According to the scale, 2.5 inches represents 5 miles. How far apart are the houses?
Found 2 solutions by josgarithmetic, jim_thompson5910:
Answer by josgarithmetic(39617)
Answer by jim_thompson5910(35256)


(distance on paper map)/(distance in real life) = (distance on paper map)/(distance in real life)

(2.5 inches)/(5 miles) = (9.25 inches)/(x miles)

2.5/5 = 9.25/x

2.5x = 5*9.25

2.5x = 46.25

2.5x/2.5 = 46.25/2.5

x = 18.5

The two houses are 18.5 miles apart